A solution is prepared by dissolving 3.50g of sodium hydroxide in water and adding water until the volume of the solution is 2.50L. What are the OH- and H3O+ concentrations?Write your answer using OH- and H3O+ identifying which you are calculating reporting three significant figures.
1
Expert's answer
2020-04-27T01:45:47-0400

Solution:

1) Molar mass of sodium hydroxide (NaOH) is 40 g/mol.

Molar mass is the mass of one mole of a substance. Therefore,

Moles of NaOH = Mass of NaOH / Molar mass of NaOH.

n(NaOH) = (3.50 g) / (40 g/mol) = 0.0875 mol.

n(NaOH) = 0.0875 mol.


2) Molarity is the number of moles of solute per liter of solution. Therefore,

Molarity of NaOH = Moles of NaOH / Volume of solution.

C(NaOH) = (0.0875 mol) / (2.50 L) = 0.0350 M.

C(NaOH) = 0.035 M.


3) NaOH is a highly soluble salt and will fully dissociate in aqueous solution to produce equivalent molarity of sodium cation and hydroxide anion:

NaOH(s) → Na+(aq) + OH(aq)

C(NaOH) = [Na+] = [OH] = 0.0350 M.

[OH] = 0.0350 M.

The concentration of OH is 0.0350 M.


4) The concentrations of hydronium (H3O+) and hydroxide (OH) ions are related together through the auto-ionization equlibrium constant (Kw):

Kw = [H3O+] × [OH] = 10-14

Therefore,

[H3O+] = Kw / [OH];

[H3O+] = (10-14) / (0.035) = 2.857×10-13 M = 2.86×10-13 M.

[H3O+] = 2.86×10-13 M.

The concentration of H3O+ is 2.86×10-13 M.


Answer:

[OH] = 0.0350 M;

[H3O+] = 2.86×10-13 M.